OpenStudy (anonymous):

simplify (-sinx-sin^2x-cos^2x)/((1+sinx)^2)

OpenStudy (anonymous):

ok i know that sin^2x+cos^2x = 1 so then -1*sinx+1/((1+sinx)^2) but where can i go from here

OpenStudy (anonymous):

(-1-sinx)/(1+sinx)^2 take - common from numerator i.e -1(1+sinx)/(1+sinx)^2 =-1/(1+sinx)

OpenStudy (anonymous):

okay thank bila

OpenStudy (anonymous):

-sinx-(sin^2x+cos^2x)/(1+sinx)^2 -sinx-1/(1+sinx)^2 =-1/(1+sinx)

OpenStudy (anonymous):

np